OpenStudy (anonymous):

Solve: x − 5 = −10 −5 5 −15 15

jimthompson5910 (jim_thompson5910):

what did you get for the answer?

OpenStudy (anonymous):

isolate x by moving the addig 5 to each side to cancel it on the one side and you are left with?

OpenStudy (anonymous):

x-5+5 = 5-10 means x = ? can you fill that in?

OpenStudy (anonymous):

-5?

jimthompson5910 (jim_thompson5910):

you got it, nice work

OpenStudy (anonymous):

good job